what is the mean and variance of samples of three cards are drawn random from a population of seven cards numbered 1 to 7


1
Expert's answer
2022-06-03T06:22:07-0400

We have population values 1,2,3,4,5,6,7 population size N=7 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{1+2+3+4+5+6+7}{7}=4"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{N}=\dfrac{9+4+1+0+1+4=9}{7}=4""\sigma=\sqrt{\sigma^2}=\sqrt{2}\approx1.4142"


The number of possible samples which can be drawn without replacement is "^{N}C_n=^{7}C_3=35."

Mean of sampling distribution 


"\mu_{\bar{X}}=\mu=4"


The variance of sampling distribution 


"Var(\bar{X})=\sigma^2_{\bar{X}}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"




"= \dfrac{4}{3}(\dfrac{7-3}{7-1})=\dfrac{8}{9}"

The number of possible samples which can be drawn with replacement is "N^n=7^3=343."

Mean of sampling distribution 


"\mu_{\bar{X}}=\mu=4"


The variance of sampling distribution 


"Var(\bar{X})=\sigma^2_{\bar{X}}= \dfrac{\sigma^2}{n}= \dfrac{4}{3}"